40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 104 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  11858   Mon Dec 7 11:17:51 2015 SteveUpdateVACnoisy RGA scan at day 433

 

Quote:

CC1 and CC2 are working again. Why did they start working  again ?

 

 

Attachment 1: 433dRGAscan.png
433dRGAscan.png
  11857   Mon Dec 7 11:11:25 2015 yutaroSummaryLSCround trip loss of X arm

On the day before yesterday and in this morning, I measured loss map of ETMX. I reported the method I used to change the beam spot on ETMX below.

Round trip loss was measured for 5 x 5 points. The result is below.

(unit: ppm)

455.4 +/- 21.1       437.1 +/- 21.8       482.3 +/- 21.8       461.6 +/- 22.5       507.9 +/- 20.1      
448.4 +/- 20.7       457.3 +/- 21.2       495.6 +/- 20.2       483.1 +/- 20.8       472.2 +/- 19.8      
436.9 +/- 19.3       444.6 +/- 19.7       483.0 +/- 19.5       474.9 +/- 20.9       498.3 +/- 18.7      
454.4 +/- 18.7       474.4 +/- 20.6       487.7 +/- 21.4       482.6 +/- 20.7       487.0 +/- 19.9      
443.7 +/- 18.6       469.9 +/- 20.2       482.8 +/- 18.7       480.9 +/- 19.5       486.1 +/- 19.2 

The correspondence between the loss shown above and the beam spot on ETMX is shown in the attached figure. In the figure, "up" and "right" indicate direction of shift of the beam spot when you watch it via the camera (ex. 455.4 ppm corresponds to the highest and rightest point in the view via the camera). 

This result is consistent withe previous result of 561.19 +/- 14.57 ppm ericq got with ASDC and reported in elog 10248 if the discussion I reported in 11819 is taken into account. Elog 11819 says in short that the strange behavior of ASDC could give us 60-70 ppm error.

The reason why the error is larger than that of the measurement for ETMY is that the noise of POX is larger than that of POY. But I am not sure to what extent the statistical error needs to be reduced.

How I shifted the beam spot on ETMX:   

Basically, the method was same as one used for Y arm. Different point is: for Y arm we have two steering mirrors TT1&2, but for X arm we have only one steering mirror BS. Then in order to shift incident beam so that the beam spot on ITMX does not change, I ran the dithering of X arm as well as that of Y arm and added offsets to both dither loops that caused same amount of shift on ETMX and ETMX. Thanks to the symmetry between X arm and Y arm, the dithering of Y arm ensured that the beam spot on ITMX was unchanged as well as that of ITMY. The idea of this method is schematically shown in Attachment 2. 

The calibration of how much the beam spot shifted is based on the results of elog 11846 . The offset was [-15,-7.5,0,7.5,15]x[-5,-2.5,0,2.5,5] for pitch and yaw, respectively.  

 

Attachment 1: image1-2.JPG
image1-2.JPG
Attachment 2: symmetry.png
symmetry.png
  11856   Mon Dec 7 10:45:21 2015 ericqUpdateCDSdaqd is mad

Since removing c1pem from the daqd master file, daqd has not crashed. I suppose we're running into the stability issue that motivated us to disable some of the other models (IOPs, RFM, etc.) during the RCG upgrade. 

A question to Jamie: although the new framebuilder prototype still had the same problem with trend writing, can it handle this higher testpoint/DQ channel load?

  11855   Mon Dec 7 10:40:09 2015 yutaroUpdateComputer Scripts / ProgramsAdded 1 line to UNFREEZE_DITHER.py

I added 1 line to one of the ASS scripts, UNFREEZE_DITHER.py like this:

L29>   ez.cawrite('C1:ASS-'+dof+'_GAIN', 0)   

The reason why I added this is: without this line, C1:ASS-'+dof+'_GAIN become larger that 1.0, which is nomial value, if you UNFREEZE DITHER when the dither is already running or C1:ASS-'+dof+'_GAIN is not 0.0.  

  11854   Mon Dec 7 00:45:28 2015 ericqUpdateCDSdaqd is mad

I glanced at the summary pages and noticed that, since Friday around when we first loaded up the new BLRMS parts, daqd has crashing very frequently (few times per hour). 

I'm going to comment out the c1pem lines from the daqd master file for tonight, and see if that helps. 

  11853   Sat Dec 5 21:57:32 2015 yutaroHowToCamerasWhen image capture does not work...

I don't know if just running "modprobe" will work or not, because I didn't try it...  When the same problem happens again, we can try just running "modprobe" first.

  11852   Sat Dec 5 21:28:33 2015 KojiHowToCamerasWhen image capture does not work...

Do we need "make" everytime? Do you mean just running "modprobe" didn't work?

  11851   Sat Dec 5 12:02:25 2015 yutaroHowToCamerasWhen image capture does not work...

Today image capture did not work again, though it had worked 3 days before. I also found that red indicator light on the front pannel of SENSORAY was not turned on, which had been turned on 3 days before (you can find SENSORAY on the floor near Pianosa). Possible reason that it did not work again was I restarted Pianosa last night. Anyway, it works now. Here I report what I did to make it work.

 

I ran thes commands in shell, following the instruction of the manual of SENSORAY 2253 (Page 5; link or you can find the manual in /users/sensoray; I put it there). 

> cd /users/sensoray/sdk_2253_1.2.2_linux

> make all

> sudo make install

> modprobe s2253

Then the red light got turned on, and image capture worked.

 

If you recieve an error like "No such file or directory: /dev/video0" at the beginning of the error message when you run image capture scripts from the medm screen, or if you notice that the red indicator light of SENSORAY is not on, this procedure could help you. 

  11850   Fri Dec 4 23:02:13 2015 yutaroUpdateLSCBeam on POX11 is possibly not centered well

[yutaro, Koji]

Now, the beam on POX11 PD is well centered and well focused.

We found out why POXDC had behaved as reported in elog 11839. There were a few reasons: the beam was not focused enough, hight of a mirror was not matched to the beam well, path of the light reflected by misaligned SRM was occasionally close to the path of POX beam.

Then, What we did is following:

- changed orientation of SRM slightly

- changed the hight of the mirror whose hight had not matched well, by changing the pedestal (hight of which mirror was changed is shown in Attachment 1.)

- put a lens with f=250 mm (where the lens is located is shown in Attachment 1.)

- refined alignment for the POX beam to hit on the center of POX11 PD.  

As a result, POX DC level behaved as shown in Attachment 2&3 when the orientation of ITMX was varied (Attachment 2: POX DC vs ITMX PIT, Attachment 3: POX DC vs ITMX YAW). 

You can see broad plateau when varied in both PIT and YAW directions, and the beam is at the center of the plateau if ITMX is aligned ideally.

 

 

Attachment 1: image1.JPG
image1.JPG
Attachment 2: 56.png
56.png
Attachment 3: 04.png
04.png
  11849   Fri Dec 4 18:20:36 2015 gautamUpdateCDSBLRMS for IMC setup

[ericq, gautam]

BLRMS filters have been set up for the coil outputs and shadow sensor signals. The signals are sent to the C1PEM model from C1MCS, where I use the library block mentioned in the previous elog to put the filters in place. Some preliminary observations:

  1. The entire operation seems to be computationally quite expensive - just for the 3 IMC mirrors, the average CPU time for C1PEM increased from ~50 usec (before any changes were made to C1PEM) to ~105 usec just as a result of installing 420 filter modules with no filter coefficients loaded (3 optics x 10 channels x 14 filter banks) to ~120 usec when all the BP and LP filters for the BLRMS blocks have been loaded and turned on (Attachment #1). Eric suggested that it may be computationally more efficient to do this without using the BLRMS library part - i.e. rather than having so many filter modules, do the RMS-ing using a piece of C code that essentially just implements the same SOS filters that FOTON generates, I will work on setting this up and checking if it makes a difference. The fact that just having empty filter modules in the model almost doubled the computation time suggests that this approach may help, but I have to think about how to implement some of the automatic checks that having a filter bank in place gives us, or if these are strictly necessary at all...
  2. While restarting the C1PEM model, we noticed that some of the optics were shaking - looking at the CPU timing signals for all the models on C1SUS revealed that both the C1SUS model and the C1MCS model were geting overclocked when C1PEM was killed (see the sharp spikes in the red and green traces in Attachment #2 - the Y scale in this plot is poor and doesnt put numbers to the overclocking, I will upload a better screenshot that Eric took once I find it). The cause is unknown.
  3. Yesterday, I noticed that when C1PEM was restarted, the states of the filter bank switches were not reverted to their states in the SDF tables. They are showing up as changes, but it is unclear why we have to manually revert them. I have also not yet added the states of the newly installed filters (BPs and LPs for the MC BLRMS blocks) to the SDF tables. 

Unrelated to this work: we cleaned up the correspondence between the accelerometer numbers and channels in the C1PEM model. Also, the 3 unused ADC blocks in C1PEM (ADC0, ADC1 and ADC2) are required and cannot be removed as the ADC blocks have to be numbered sequentially and the signals needed in C1PEM come from ADC3 (as we found out when we tried recompiling the model after deleting these blocks).

Attachment 1: c1pem_timing.png
c1pem_timing.png
Attachment 2: C1SUS_overclocked.png
C1SUS_overclocked.png
  11848   Fri Dec 4 13:12:41 2015 ericqUpdateCDSc1ioo Timing Overruns Solved

I had noticed for a while that the c1ioo frontend model had much higher variability than any of the other other 16k models, and would run longer than 60us multiple times an hour. This struck me as odd, since all it does is control the WFS loops. (You can see this on the Nov17 Summary page. (But somehow, the CDS tab seems broken since then, and I'm not sure why...))

This has happily now been solved! While poking around the model, I noticed that the MC2 transmission QPD data streams being sent over from c1mcs were using RFM blocks. This seemed weird to me, since I wasn't even aware that c1ioo had an RFM card. Since the c1sus and c1ioo frontends are both on the Dolphin network, I changed them to Dolphin blocks and voila! The cylce time now holds steady at 21usec. 


Update: I think I figured out the problem with the CDS summary pages. Looking at the .err files in /home/40m/public_html/summary/logs on the 40m LDAS account showed that C1:FEC-33_CPU_METER wasn't found in the frame files. Indeed, this channel was commented out in c1/chans/daq/C0EDCU.ini. I enabled it and restarted daqd. Hopefully the CDS tab will come back soon...

  11847   Fri Dec 4 12:33:52 2015 yutaroUpdateLSCBeam on POX11 is possibly not centered well

To focus POX beam on POX11 PD, I added an iris and a lens before POX11 PD as you can see in Attachment 1.

It seemed that the beam is well focused, but the behavior of POXDC has not changed, as shown in Attachments 2 & 3.    

Attachment 1: image1-3.JPG
image1-3.JPG
Attachment 2: 07.png
07.png
Attachment 3: 47.png
47.png
  11846   Fri Dec 4 10:18:33 2015 yutaroUpdateASSOffset in the dither loop of XARM vs beam spot shift on ETMX

As I did for YARM (elog 11779), I measured the relation between offsets added just after the demodulation of the dithering loop of XARM and beam spot shift on ETMX. Defferent from YARM, the beam spot on ITMX DOES change because only BS is used as a steering mirror (TT1&2 are used for the dithering of YARM). Instead, the beam spot on BS DOES NOT change.

This time, I measured by oplevs the angles of both ETMX and ITMX for each value of offset, and using these angles I calculated the shift of the beam spot on ETMX so that I got two independent estimations (one from ETMX oplev, and the other from ITMX oplev) as shown below. The calibration of the oplevs reported in elog 11831 is taken into account. 

The difference of two estimations comes from the error of calibration of oplevs and/or imperfect alignment, I think. 

Attachment 1: offset-angleETMXPIT.png
offset-angleETMXPIT.png
Attachment 2: offset-angleETMXYAW.png
offset-angleETMXYAW.png
  11845   Thu Dec 3 19:10:28 2015 yutaroUpdateLSCXARM lock with ITMX actuated and related change on ASS

To avoid the strange kicking of ETMX, I locked XARM with ITMX actuated instead of ETMX so that I changed elements of C1LSC_OUTPUT_MTRX; before: XARM=ETMX, after: XARM=ITMX.

And I change C1:LSC-XARM_GAIN from 0.007 to 0.022.

 

With this setup, I ran dither but then error signals of dithering oscillated as shown in the figure below.

Then I found that if C1:ASS-XARM_ETM_PIT_L_DEMOD_SIG_GAIN / C1:ASS-XARM_ETM_YAW_L_DEMOD_SIG_GAIN in C1ASS_LOCKINS_XARM.adl are changed as 0.200 -> 0.100 and 0.200 -> 0.100, respectively, the dithering works well.

But the burt file that is loaded when you let dithering "ON" is not changed, because now I don't know how to update a burt file. So, if you let dithering "ON", the dithering will run with the condition that C1:ASS-XARM_ETM_PIT_L_DEMOD_SIG_GAIN / C1:ASS-XARM_ETM_YAW_L_DEMOD_SIG_GAIN are not 0.100 but 0.200.

 

   

Attachment 1: 40.png
40.png
  11844   Thu Dec 3 18:18:48 2015 gautamUpdateCDSBLRMS for optics suspensions - library block

In order to be consistent with the naming conventions for the new BLRMS filters, I made a library block that takes all the input signals of interest (i.e. for a generic optic, the coil signals, the local damping shadow sensors, and the Oplev Pitch and Yaw signals - so a total of 12 signals, unused ones can be grounded). The block is called "sus_single_BLRMS". Inside the block, I've put in 12 BLRMS library blocks, with each input signal going to one of them. All the 7 outputs of the BLRMS block are terminated (I got a compiling error if I did not do this). The idea is to identify the optic using this block, e.g. MC2_BLRMS. The BLRMS filters inside are called UL_COIL, UR_COIL etc, so the BLRMS channels will end up being called C1:SUS-MC2_BLRMS_UL_COIL_0p01_0p03 and so on. I tried implementing this in C1PEM, but immediately after compiling and restarting the model, I noticed some strange behaviour in the seismic rainbow STS strip in the control room - this was right after the model was restarted, before I attempted to make any changes to the C1PEM.txt file and add filters. I then manually opened up the filter bank screens for the RMS_STS1Z bandpass and lowpass filters, and saw that the filter switches were OFF - I wonder if this has something to do with these settings not being updated in the SDF tables? So I manually turned them on and cleared the filter hitsory for all 7 low pass and band pass filter banks, but the traces on the seismic striptool did not return to their nominal levels. I manually checked the filter shapes with Foton and they seem alright. Anyways, for now, I've reverted to the C1PEM model before I made any changes, and the seismic strip looks to be back at its normal level - when I recompiled and restarted the model with the changes I made removed, the STS1Z BLRMS bandpass and lowpass filters were ON by default again! I'm not sure what I'm doing wrong here, I will investigate this further. 

  11843   Thu Dec 3 10:05:07 2015 yutaroUpdateOptical LeversCalibration of oplevs for ITMX/ETMX

I updated the figures. I think it's easier to read now.

  11842   Thu Dec 3 06:15:38 2015 ranaUpdateOptical LeversCalibration of oplevs for ITMX/ETMX

http://blogs.mathworks.com/loren/2007/12/11/making-pretty-graphs/

Let Loren help you make your Oplev data readable to humans.cool

  11841   Thu Dec 3 03:01:07 2015 gautamUpdateLSCCalibration of C1CAL

[ericq, gautam]

While trying to resolve the strange SRCL loop shape seen yesterday (which has been resolved, eric will elog about it later), we got a chance to put in the correct filters to the "CINV" branch in the C1CAL model for MICH, PRCL, and SRCL - so we have some calibrated spectra now (Attachment #1). The procedure followed was as follows:

  1. Turn on the LO gain for the relevant channel (we used 50 for MICH and SRCL, 5 for PRCL)
  2. Look at the power spectra of the outputs of the "A" and "CINV" filter banks - the former has some calibrated filters in place already (though I believe they have not accounted for everything).
  3. Find the peak height at the LO excitation frequency for the two spectra, and calculate their ratio. Use this to install a gain filter in the CINV filter module for that channel. 
  4. Look at the spectrum of the output of the "W" filter bank for that channel - the plot attached shows this information.

The final set of gains used were:

MICH: -247 dB

PRCL: -256 dB

SRCL: -212 dB

and the gain-only filters in the CINV filter banks are all called "DRMI1f".

Once we are able to lock the DRFPMI again, we can do the same for CARM and DARM as well...

Attachment 1: 2015-12-C1CAL_Calibration.pdf
2015-12-C1CAL_Calibration.pdf
  11840   Wed Dec 2 18:54:20 2015 gautamUpdateCDSChanges to C1MCS, C1PEM

Summary:

I've made several changes to the C1MCS model and C1PEM model, and have installed BLRMS filters for the MC mirror coils, which are now running. The main idea behind this test was to see how much CPU time was added as a result of setting up IPC channels to take the signals from C1MCS to C1PEM (where the BLRMS filtering happens) - I checked the average CPU time before and after installing the BLRMS filters, and saw that the increase was about 1 usec for 15 IPC channels installed (it increased from ~27usec to 28usec). A direct scaling would suggest that setting up the BLRMS for the vertex optics might push the c1sus model close to timing out - it is at ~50usec right now, and I would need, per optic, 12 IPC channels, and so for the 5 vertex optics, this would suggest that the CPU timing would be ~55usecI have not committed either of the changed models to the SVN just yet

Details:

  1. On Eric's suggestion, I edited the C1_SUS_SINGLE_CONTROL library part to tap the signal at the input to the output filters to the coils, as this is what we want the BLRMS of. I essentially added 5 more outputs to this part, one for each of the coils, and they are named ULIn etc to differentiate it from the signal after the output filters. I have not committed the changed library part to the SVN just yet
  2. I used the cdsIPCx_SHMEM part to pipe the signals from C1MCS to C1PEM - a total of 15 such channels were required for the three IMC mirrors.
  3. I added the same cdsIPCx_SHMEM parts to the C1PEM model in the receiving configuration, and connected their outputs to BLRMS blocks. The BLRMS blocks themselves are named as RMS_MC2_UL_COIL_IN and so on.
  4. I shutdown the watchdogs to MC1, MC2 and MC3, and restarted the C1PEM and C1MCS models on C1SUS. Yutaro pointed out that on restarting C1MCS, the IMC autolocker was disabled by default - I have enabled it again manually.
  5. I was under the impression that each time a BLRMS block is added, a filter bank is automatically added to the C1PEM.txt file in /opt/rtcds/caltech/c1/chans - turns out, it doesn't and my script for copying the template bandpass and lowpass filtes into the .txt files was needlessly complicated. It suffices to change the filter names in the template file, and append the template file to C1PEM.txt using the cat command: i.e. cat template.txt > C1PEM.txt. The computer generated file seems to organize the filters in alphabetical order, and my approach obviously does not do so, but the coefficients are loaded correctlty and the filters seem to be functioning correctly so I don't think this is a problem (I measured the transfer function of one of the filters with DTT, it seems to match up well with the Foton bode plot). 
  6. I added a few lines to the script to also turn on the filter switches after loading the filter coefficients.

Next steps:

Now that I have a procedure in place to install the BLRMS filters, we can do so for other channels as well, such as for the coils and Oplevs of the vertex optics, and the remaining PEM channels (SEIS, accelerometers, microphones?). For the vertex optics though, I am not sure if we need to do some rearrangement to the c1sus model to make sure it does not time out...

  11839   Wed Dec 2 17:14:33 2015 yutaroUpdateLSCBeam on POX11 is possibly not centered well

I checked how POXDC level changes when the angle of ITMX is varied. ETMX was misaligned.

Then I found that in YAW direction the POXDC level is maximized but it doesnt have plateau, and in PIT direction it is not maximized so that it is at the slope and it doesnt have plateau, as shown in attached figures. These results indicate that the beam size on POX11 is not small enough compared to the size of the diode and it is not centered well.

Attachment 1: 47.png
47.png
Attachment 2: 41.png
41.png
  11838   Wed Dec 2 15:52:28 2015 yutaroUpdateLSCrestoration of POXDC

I disconnected the cable that was connected to CH6 of the whitening filter in 1Y2, then connected POXDC cable to there (CH6). This channel is where POXDC used to connect.

Then I turned on the whitening filter for POXDC and POYDC (C1:LSC-POXDC FM1, C1:LSC-POYDC FM1) and changed the gain of analog whitening filter for POXDC and POYDC from 0 dB to 45 dB and from 0 dB to 39 dB, respectively (C1:LSC-POXDC_WhiteGain, C1:LSC-POYDC_WhiteGain).

  11837   Wed Dec 2 15:08:41 2015 ericqUpdateComputer Scripts / ProgramsDonatella sudo problem resolved

Somehow, the controls user account on donatella lost its membership to the sudoers group, which meant doing anything that needs root authentication was impossible. 

I fixed this by booting up from a Linux install USB drive, mounting the HD, and running useradd controls sudo

  11836   Wed Dec 2 03:34:30 2015 ericqUpdateLSCSRCL OLG weirdness

[gautam, ericq]

Since ETMX seems to have been on good behavior lately, we tried to fire the IFO back up. 

We had a fair amount of trouble locking the DRMI with the arms held off resonance. For reasons yet to be understood, we discovered that the SRCL OLG looks totally bananas. It isn't possible to hold the DRMI for very long with this shape, obviously. 

With the arms misaligned and the DRMI locked on 1F, the loop shape is totally normal. I haven't yet tried 3F locking with the arms misaligned, but this is a logical next step; I just need to look up the old demod angles used for this, since it wasn't quickly possible with the 3F demod angles that are currently set for the DRFPMI. 

Attachment 1: weirdSRCLG.pdf
weirdSRCLG.pdf
  11835   Tue Dec 1 20:20:16 2015 KojiUpdateGeneralMegatron maitenence

MC Autolocker got stack somewhere. I had to go to megatron and kill MC Autolocker.

init relaunched the autolocker automatically, and now it started properly.

  11834   Tue Dec 1 17:26:14 2015 KojiUpdateGeneralMegatron maitenence

SLOWDC servo was dead. I followed EricQ's instruction.

  11833   Tue Dec 1 17:16:31 2015 gautamUpdateCDSScript for copying BLRMS filters

We've been talking about putting in BLRMS filters for several channels - it would be a pain to manually copy over the correct bandpass and lowpass filter coefficients into the newly created filter banks, and so I've set up a script (attached) that can do the job. As template filters, I'vm using the filters rana detailed here. Essentially, what the script does is identify the (empty on creation) block of text for a given filter: e.g. RMS_STS1Z_BP_0p01_0p03 for STS1Z), and appends the template filter coefficients. To test my script, I first backed up the original C1PEM.txt file from /opt/rtcds/caltech/c1/chans, removed all the filter coefficients for the STS1Z BLRMS filters, and then replaced it with one generated using my script. I then loaded the coefficients for all the filters in the C1PEM modules, without any obvious error messages being generated. I also checked that foton could read the new file, and checked tmake sure that sensible filter shapes were seen for some channels. Since this seems to be working, I'm going to start putting in BLRMS blocks into the models tomorrow.

Attachment 1: makeFilterFiles.bash.zip
  11832   Tue Dec 1 16:55:04 2015 SteveUpdateVACDataviewer fixed

Q adjusted the Dataviewer so it is not chopping of data any more. Thanks.

 

Cold cathode gauge reading of 10 years.

 

Attachment 1: 10yrsCC1&4M.png
10yrsCC1&4M.png
  11831   Tue Dec 1 11:26:23 2015 yutaroUpdateOptical LeversCalibration of oplevs for ITMX/ETMX

With the same method as reported in elog 11785, I calibrated oplevs for ITMX/ETMX.

 

RESULTS 

 

According to this measurement, ratio of the calibration factor derived with this measurement (NEW) and the calibration factor for now (OLD), i.e. NEW/OLD was:   

ETMX_PIT: 4.470

ETMX_YAW: 2.5970

ITMX_PIT: (-)1.1102

ITMX_YAW: 1.8173

 

NOTE

The calibration factors of the oplevs for ETMY/ITMY are   NOT UPDATED YET. I updated on Dec 11, 2015

 

Attachment 1: ep_l.pdf
ep_l.pdf
Attachment 2: ey_l.pdf
ey_l.pdf
Attachment 3: ip_l.pdf
ip_l.pdf
Attachment 4: iy_l.pdf
iy_l.pdf
  11830   Tue Dec 1 10:52:52 2015 SteveUpdateGeneralDataviewer

Dataviewer x axis end is not there.

On ( 2600 days) longer plots it is missing 8 moths and on (100 days) shorther plot it is missing 1 month.

Attachment 1: xAxisEndMissing.png
xAxisEndMissing.png
  11829   Mon Nov 30 18:27:30 2015 KojiUpdateLSCstrange behavior of ASDC

It wasn't fully mentioned in ELOG 11814.
We checked the PD first and this behavior didn't change after the realignment of the AS55PD.
Yutaro confirmed that this effect is happening in the vacuum chamber.

  11828   Mon Nov 30 17:17:30 2015 yutaroUpdateLSCDoes a baffle in front of ETMY have effect on loss map measurement?

With captured images of ETMYF, I measured the shift of the beam spot on ETMY.

The conclusions are:

the baffle would have almost no effect on loss map measurement and

the calibration of beam spot shift is confirmed to be not so bad.

 

What I did:

I captured ETMYF images in the cases that (i)beam spot is centered on ETMY, beam spot is at the rightest and lowest point of my loss map measurement (corresponding to [0,0] component of the matrix shown in elog 11818), and beam spot is at the leftest and highest point of my loss map measurement ([4,4] component). Each captured image is attached.

Then using ImageJ, I measured the shift of the beam spot. I calibrated lengh in horizontal direction and vertical direction with the diameter of the mirror. 

Results:

The amount of the beam shift was 7.2 mm and 8.0 mm for each case.

These values indicate that clapping loss due to the baffle is less than 10 ppm in a round trip.

Today's results support the previous calibration with oplev, which says the amount of the beam shift is 7.0 mm. Two values derived by different calibrations coincide within ~10 % though they are totally different methods. This also support the calibration of the oplev for ETMY (elog 11785) indirectly. 

 

 

Attachment 1: element2-2.png
element2-2.png
Attachment 2: element0-0.png
element0-0.png
Attachment 3: element4-4.png
element4-4.png
  11827   Mon Nov 30 16:33:06 2015 ericqUpdateLSCstrange behavior of ASDC

One possible explanation of this behavior is simply poor centering of the AS beam on AS55 (whose DC level provides ASDC, if memory serves me correctly).

I misaligned ETMY, and moved ITMY through its current nominal alignment while looking at the POYDC and ASDC levels. 

In both pitch and yaw, the nominal alignment is fairly close to the "plateau" in which the AS beam is fully within the PD active surface. I.e. it doesn't take much angular motion to start to lose part of the beam, and thus introduce a first order coupling of angle to power. (Look at the plateaus at around -2min and -0.5min, and where the rapidly changing oplev trace crosses zero)

Furthermore, POYDC seems to be in some weird condition where it is actually possible to increase the reported powerwhen misaligning in pitch, but somehow there is more angular coupling in this state. 

In any case, I would advise that the POY11 and AS55 RFPDs have their spots recentered with optics in their nominal aligned states. In fact, given how we found REFL11 alingment to be less-than-ideal not so long ago, all of the RFPDs could probably use a checkup. 

  11826   Mon Nov 30 15:17:57 2015 KojiUpdateLSCLO level check for the LSC RF distribution box

T1000461 tells us that the nominal LO input is 2dBm although we don't know what's the LO level is at the mixers in the demod boards.

  11825   Mon Nov 30 14:12:14 2015 ericqUpdateLSCLO level check for the LSC RF distribution box

I checked the RF levels at the LSC LO distribution box, with the agilent scope and a handful of couplers. This was all done with the Marconi at +13dBm. 

I only checked the channels that are currently in use, since the analyzer only measures 3 channels at a time, and rewiring involves walking back and forth to the IOO rack to make sure unpowered amps aren't driven, and I was getting hungry. 

For the most part, the LO levels coming into the LSC demod boards are all around +1.5dBm (i.e. I measured around -18.0dBm out of the ZFDC-20-5 coupler, which has a nominal 19.5dB coupling factor)

The inputs piped over from the IOO rack, labeled as "+6dBm" were found to be 4.7dBm and 2.9dBm for 11Mhz and 55MHz, respectively. 

The 2F signals were generally about 40dB lower, with two exceptions:

  • REFL165's ~332MHz signal was around -18dBc
  • POP22 had many more visible harmonics than any other LO signal
    • 11MHz: -32 dBc
    • 33MHz: -32 dBc
    • 44Mhz: -15dBc 

Here are the raw numbers I measured out of the couplers, all in dBm:

  • 11MHz in: -14.8
  • 55MHz in: -16.6
  • POX11:    -18.7
  • POY11:    -18.0
  • REFL11:   -18.0
  • REFL33:   -18.3
  • POP110:   -17.9
  • AS110:    -18.1
  • POP22:    -19.9
  • REFL165:  -18.5
  • AS55:     -18.6
  • POP55:    -18.8 (this port is used as the REFL55 LO)
  11824   Mon Nov 30 12:19:38 2015 yutaroUpdateComputer Scripts / Programsimage capture

On VIDEO.adl, Image Capture and Video Capture did not seem to work and gave me some errors, so I fixed following two things:

1. just put one side of a USB cable to Pianosa the other side of which was connected to Sensoray; I don't know why but this was unconnected.

2. slightly fixed /users/sensoray/sdk_2253_1.2.2_linux/imsub/display-image.py as fpllows   

L52:       pix[j, i] = R, G, B     ->     pix[j, i] = int(R), int(G), int(B)  

It seems to work, at least for some cameras including ETMYF and ITMYF. 

  11823   Mon Nov 30 10:41:45 2015 yutaroUpdateLSCDoes a baffle in front of ETMY have effect on loss map measurement?

It might have, so I think I need to estimate shift of beam spot more preciely.

 

According to Steve's drawing, radius of the hole of the baffle is 19.8 mm.

Intensity distribution of fundamental mode in x axis direction is this (y is integrated out):

I(x)=\frac{1}{\sqrt{2\pi(w/2)^2}}e^{-\frac{1}{2}\left(\frac{x}{w/2}\right)^2}

With the radius of curvature of ETMY of 60 m and the arm length of 37.78 m, the beam width w on ETMY is estimated to be 5.14 mm. From this expression of the intensity, \int^\infty_{x=9.56\mathrm{mm}}\mathrm{d}xI(x)=100\,\mathrm{ppm},\,\int^\infty_{x=10.00\mathrm{mm}}\mathrm{d}xI(x)=50\,\mathrm{ppm}, for example. If round trip loss is considered, these values are doubled.

Although maximum shift of beam spot from the ideal spot on ETMY is estimated to be sqrt(6.0^2+(1.7+1.7)^2)=6.9 mm, this value could have error of several tens of % because I am not sure to what exten the calibration is precise, which means that the maximum shift could be ~10 mm and seperation between the baffle and the beam could be ~10 mm.

Therefore, I need to check how much the beam spot shifts with another way, maybe with captured image of the CCD camera. 

 

  11822   Sun Nov 29 12:32:26 2015 KojiUpdateLSCCurrent state of the frequency source, and possible improvement

I'm not claiming we need to modify the frequency source immediately as we are not limited by the oscillator amplitude or phase noise.
I just wanted to note something in mind before it goes away quickly.

Alberto's T1000461 tells us that the oscillator and phase noise are degraded by factor of ~3 and ~5 due to the RF chanin.
My diagram is possible removal of up-down situation of the chain.

Maybe more direct improvement would be:

- Removal of two amplifiers out of four. The heat condition of the box is touch thought it is not critical.

- The modification will allow us to have a spare 11MHz channel at 1X2 rack that would be useful for 3f modulation.

  11821   Sun Nov 29 05:23:57 2015 ranaUpdateLSCCurrent state of the frequency source, and possible improvement

I need some more hints to understand the improvement, although its generally good to re-build it considering the sad state of the assembly/installation that you found.

I see that the current design brings the 11 MHz signal to -2 dBm before intering the first ZHL-2+, but since that has a NF of 9 dB, that seems to only degrade the phase noise to -2 - (-174 +9) = -163 dBc. That seems OK since we only need -160 dBc from this system. Probably the AM noise is worse than this already (we should remember to hook up a simple AM stabilizer in 2016, as well as the ISS).

What else are the main features of this improvement? I can reward a good summary with some Wagonga.

  11820   Sat Nov 28 11:46:40 2015 yutaroUpdateLSCpossible error source of loss map measurement

I found that TRY level degraded and the beam shape seen with CCD camera at AS port was splitted when the beam spot on ETMY was not close to the center. This was because dither started not working well. I suspect so because in such a case TRY level went up when I did iteration with TT1 and TT2 after freezing dither. Splitted beam shape indicates that incident light did not match well with the cavity mode.

TRY level for each point was this:

TRYDC
[[ 0.6573      0.8301      0.8983      0.8684      0.6773    ]
 [ 0.7555      0.8904      0.9394      0.8521      0.6779    ]
 [ 0.6844      0.8438      0.9318      0.8834      0.6593    ]
 [ 0.7429      0.8688      0.9254      0.8427      0.6474    ]
 [ 0.7034      0.8447      0.8834      0.8147      0.6966    ]]

 In the worst case, TRY level was 70 % of the maximum level. Assuming that this degrade was totally due to the mode mismatch, this corresponds to ~50 urad difference between the angle of incident light and resonant lighe in the arm (see elog 11819).

  11819   Fri Nov 27 22:20:24 2015 yutaroUpdateLSCround trip loss of Y arm

Here, I upload data I took last night, including the power of reflected power (locked/misaligned) and transmitted power for each point (attachement 1).

 

And I would like to write about possible reason why the loss I measured with POYDC and the loss I measured with ASDC are different by about 60 - 70 ppm (elog 11810 and 11818). The conclusion I have reached is: 

It could be due to the strange bahavior of ASDC level. 

This difference corresponds to the error of ~2% in the value of P_L/P_M. As reported in elog 11815, ASDC level changes when angle of the light reflected by ITMY changes, and 2% change of ASDC level corresponds to 10 urad change of the angle of the light according to my rough estimation with the figure shown in elog 11815 and attachment 2. This means that 2% error in P_L/P_M could occur if the angle of the light incident to YARM and that of resonant light in YARM differ by 10 urad. Since the waist width w_0 of the beam is ~3 mm, with the 10 urad difference, the ratio of the power of TEM10 mode is (10\,\mu \mathrm{rad}/ \theta_0)^2\sim0.01, where \theta_0=\lambda/\pi w_0. This value is reasonable; in elog 11743 Gautam reported that the ratio of the power of TEM10 was ~ 0.03, from the result of cavity scan. Therefore it is possible that the angle of the light incident to YARM and that of resonant light in YARM differ by 10 urad and this difference causes the error of ~2% in P_L/P_M, which could exlain the 60 - 70 ppm difference. 

Attachment 1: log.txt.zip
Attachment 2: 17.png
17.png
  11818   Fri Nov 27 03:38:23 2015 yutaroSummaryLSCround trip loss of Y arm

Tonight I measured "loss map" of ETMY. The method to calculate round trip loss is same as written in elog 11810, except that I used POYDC instead of ASDC this time.

How I changed beam spot on ETMY is: elog 11779.

I measured round trip loss for 5 x 5 points. The result is below.

(unit: ppm)

494.9 +/- 7.6       356.8 +/- 6.0       253.9 +/- 7.9       250.3 +/- 8.2       290.6 +/- 5.1      
215.7 +/- 4.8       225.6 +/- 5.7       235.1 +/- 7.0       284.4 +/- 5.4       294.7 +/- 4.5      
205.2 +/- 6.0       227.9 +/- 5.8       229.4 +/- 7.2       280.5 +/- 6.3       320.9 +/- 4.3      
227.9 +/- 5.7       230.5 +/- 5.5       262.1 +/- 5.9       315.3 +/- 4.7       346.8 +/- 4.2      
239.7 +/- 4.5       260.7 +/- 5.3       281.2 +/- 5.8       333.7 +/- 5.0       373.8 +/- 4.9 

The correspondence between the loss shown above and the beam spot on ETMY is shown in the following figure. In the figure, "downward" and "left" indicate direction of shift of the beam spot when you watch it via the camera (ex. 494.9 ppm corresponds to the lowest and rightest point). 

Edited below on 28th Nov. 

To shift the beam spot on ETMY, I added offset in YARM dither loop. The offset was [-30,-15,0,15,30]x[-10,-5,0,5,10] for pitch and yaw, respectively.  How I calibrated the beam spot is basically based on elog 11779, but I multiplied 5.3922 for vertical direction and 4.6205 for horizontal direction which I had obtained by caliblation of oplev (elog 11785).

Edited above on 28 th Nov.    

 

   

 

I will report the detail later. 

  11817   Thu Nov 26 19:39:27 2015 KojiUpdateLSCCurrent state of the frequency source, and possible improvement

Uploaded on T1000461 too.

Attachment 1: RF_Frequency_Source.pdf
RF_Frequency_Source.pdf RF_Frequency_Source.pdf
  11816   Wed Nov 25 23:34:52 2015 yutaroUpdateLSCround trip loss of Y arm

[yutaro, Koji]

Due to the strange behavior (elog 11815) of ASDC level, we checked if it is possible to use POYDC instead of ASDC to measure the power of reflected light of YARM. Attached below is the spectrum of them when the arm is locked. This spectrum shows that it is not bad to use POYDC, in terms of noise. The spectrum of them when ETMY is misaligned looked similar.

So I am going to use POYDC instead of ASDC to measure arm loss of YARM. 


Ed by KA:
The spectra of POYDC and ASDC were measured. We foudn that they have coherence at around 1Hz (good).
It told us that POYDC is about 1/50 smaller than ASDC. Therefore in the attached plot, POYDC x50 is shown.
That's the meaning of the vertical axis unit "ASDC".

Attachment 1: 14.png
14.png
  11815   Wed Nov 25 23:17:34 2015 yutaroUpdateLSCstrange behavior of ASDC

[yutaro, Koji]

I noticed that ASDC level changes depending on the angle of ITMY when trying to take some data for loss map of YARM. We finally found that ASDC level behaves strangely when the angle of ITMY in yaw direction is varied, as you can see in Attachment 1. Now, AS port recieved only the reflection of ITMY. 

NOTE: This behavior indicates that angular motion could couple to length signal in AS port.      

Koji suggested that this behavior might be caused by interference at SR2 or SR3 between main path light and the light reflected by the AR surface. By rough estimation, we confirmed that this scenario would be possible. So it would be better to measure AR reflection of the same mirror to ones used for SR2 and SR3 in term of incident angle.     

Ed by KA: This senario could be true if the AR reflection of teh G&H mirrors have several % due to large angle of incidence. But then we still need think about the overlap between the ghost beam and the main beam. It's not so trivial.

Attachment 1: 14.png
14.png
  11814   Wed Nov 25 22:59:42 2015 yutaroUpdateLSCAS table optics realignment

I slightly changed the orientation of a few mirrors on AS table that are used to make the AS light get into PDs, in order to confirm that the strange behavior of ASDC (I will report later) is not caused by clipping related to these mirrors or miscentering on PDs.

Then output level of ASDC, AS55, and AS165 could have changed.   

So take care of this possible change when you do something related to them. But the relative change of them would be at most several %, I think.

 

  11813   Wed Nov 25 22:37:12 2015 yutaroUpdateLSCremoval of Gautam's cable in 1Y2 and restoration of POYDC

[yutaro, Koji]

We disconnected the cable that was connected to CH5 of the whitening filter in 1Y2, then connected POYDC cable to there (CH5). This channel is where POYDC used to connect.

Then we turned on the whitening filter for POYDC (C1:LSC-POYDC FM1) and changed the gain of analog whitening filter for POYDC from 0 dB to 39 dB (C1:LSC-POYDC_WhiteGain).

  11812   Wed Nov 25 20:07:35 2015 ranaUpdateIOOLO level check for the IMC demod board

Perhaps we can replace T1 with a mini-circuits hybrid 0-90 deg splitter and then remove the trim caps. (JSPQ-80, JYPQ-30, SCPQ-50)

  11811   Wed Nov 25 16:46:10 2015 KojiUpdateIOOLO level check for the IMC demod board

Awwww. I found that the demod board has the power splitter (PSCJ-2-1) with one output unterminated.
This power splitter should be removed.

https://dcc.ligo.org/D1500443

Attachment 1: D990511-40m_151123.pdf
D990511-40m_151123.pdf
  11810   Wed Nov 25 16:40:32 2015 yutaroUpdateLSCround trip loss of Y arm

I measured round trip loss of Y arm. The alignment of relevant mirrors was set ideal with dithering (no offset).

Summary:

 round trip loss of Y arm: 166.2 +/- 9.3 ppm

(In the error, only statistic error is included.)

How I measured it: 

I compared the power of light reflected by Y arm (measured at AS) when the arm was locked (P_L) and when ETMY was misaligned (P_M). P_L and P_M can be described as 

P_M=P_0(1-T_\mathrm{ITM})

P_L=P_0\left[1-(1-\alpha)\frac{4T_\mathrm{ITM}}{T_\mathrm{tot}^2}T_\mathrm{loss}\right].

The reason why P_L takes this form is: (1-alpha)*4T_ITM/(T_tot)^2 is intracavity power and then product of intracavity power and loss describes the power of light that is not reflected back. Here, alpha is power ratio of light that does not resonate in the arm (power of mismatched mode and modulated sideband), and T_tot is T_ITM+T_loss. Transmissivity of ETM is included in T_loss. I assumed alpha = 7%(mode mismatch) + 2 % (modulation) (elog 11745)

After some calculation we get

1-P_L/P_M\simeq \frac{4(1-\alpha) T_\mathrm{loss}}{T_\mathrm{ITM}}-T_\mathrm{ITM}.

Here, higher order terms of T_ITM and (T_loss/T_ITM) are ignored. Then we get

(1-\alpha) T_\mathrm{loss}=\frac{T_\mathrm{ITM}}{4}(1-P_L/P_M+T_\mathrm{ITM}).

Using this formula, I calculated T_loss. P_L and P_M were measured 100 times (each measurement consisted of 1.5 sec ave.) each and I took average of them. T_ETM =13.7 ppm is used.

Discussion: 

-- This value is not so different from the value ericq reported in July (elog 10248).

-- This method of measuring arm loss is NOT sensitive to T_ITM.  In contrast, the method in which loss is obtained from finesse (for example, elog 11740) is sensitive to T_ITM.

In the method I'm now reporting, 

\Delta T_\mathrm{loss}/T_\mathrm{loss}\simeq\Delta T_\mathrm{ITM}/T_\mathrm{ITM},

but in the method with finesse,

\Delta T_\mathrm{loss}\simeq\Delta T_\mathrm{ITM}.

In the latter case, if relative error of T_ITM is 10%, error of T_loss would be 1000 ppm.

So it would be better to use power of reflected light when you want to measure arm loss.  

  11809   Wed Nov 25 14:46:53 2015 gautamUpdateCDSc1lsc models restarted

I noticed that all the models running on C1LSC had crashed when I came in earlier today. I restarted all of them by ssh-ing into C1LSC and running rtcds restart all. The models seem to be running fine now.

ELOG V3.1.3-